OpenStudy (anonymous):

find the slope of (2,-2) and (0,-1)

OpenStudy (agreene):

\[\frac{\Delta y}{\Delta x}\] \[=\frac{2-0}{-2-(-1)}=\frac{2}{-2+1}=\frac{2}{-1}=-2\]

OpenStudy (anonymous):

oh. i got \[-2/1\] and didn't symplify it into -2

OpenStudy (anonymous):

that is still wrong. it doesn't fit

OpenStudy (anonymous):

its y2-y1 over x2-x1

OpenStudy (anonymous):

Slope is the change in y's over the change in x's. -1 - -2 -1 + 2 1 ------- = ------ = -- = -1/2 0 - 2 -2 -2

OpenStudy (anonymous):

i have been doing xs over ys i think

OpenStudy (agreene):

lol I messed mine up something awesome... i even said to do y/x and then i did x/y... my bad

OpenStudy (anonymous):

so -2 is wrong???

OpenStudy (anonymous):

okay. now i got 0/2

OpenStudy (agreene):

\[\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}=\frac{-2-(-1)}{2-0}=\frac{-2+1}{2}=-\frac{1}{2}\]

OpenStudy (anonymous):

where did the one come from

OpenStudy (anonymous):

a negative and a negative is a positive

OpenStudy (anonymous):

oh my gosh. i got it. nevermind. sorry.

OpenStudy (anonymous):

haha! -_-